Manchester Conditioning Update 2

Sophie’s New Choice: Temperature, Flow or pH?

This week, I’ve been at Manchester trying to balance temperature, flow and pH in the conditioning set-up. On June 25, Laura told me that the pH in my tanks was reading very low (around 7.3). With all of the excess feeding, the oysters are respiring so much that they’re modifying the chemistry in the tanks.

Monday, June 26

I went ot Manchester to adjust the flow rate and raise the pH. First, I vacuumed the tanks and checked mortality. One oyster died recently from Tank 6A. I increased the flow to 1.5 L/min from 1.2 L/min, which raised the pH to 7.7. I then fed the oysters 490 mL of Reed’s Shellfish Diet in 200 L of freshwater, dosing at 55%. When I left, the pH was between 7.65 and 7.7.

Wednesday, June 28

After monitoring pH for a day, I found it was still dropping below 7.6 in my conditioning tanks. Additionally, the 800 W heater was unable to maintain my desired temperature with the increased flow. I drained and cleaned the oyster tanks, finding one dead oyster in Heat Shock A. I increased the flow to 2.4 L/min which stabilized the pH once more, but caused temperature to drop.

Thursday, June 29

First, I vacuumed the oyster tanks and bleached the algal line. No mortalities found!

Tag Label A B Total
1 8 9 17
2 9 8 17
3 7 6 13
4 9 7 16
5 9 8 17
6 7 8 15
Heat Shock 5 6 11
Full Amb 7 8 15
Spare 2 3 5
Total 63 63 126

Table 1. Current oyster count in conditioning set-up.

While feeding the oysters, I stabilized the pH at 7.7-7.8 by increasing flow to 2.7 L/min.


Figure 1. pH and temperature reading after flow stabilization.

I took another 800 W heater from the lab and added it to my header tank. I set the original heater at 16 ºC. The second heating unit’s control box was in Farenheit, so I set it to 68 ºF to see what would happen.


Figure 2. In order to plug in the second heater, Olivia and I had to move the AVTECH unit to a different outlet.



Figures 3-4. Temperature settings on controllers. I wanted to the heated water temperature to be 16 ºC, so I set the old heater setpoint to 16 ºC and the new one to 68 ºF.

screen shot 2017-06-30 at 10 53 02 am

Figure 5. AVTECH computed pH values over the past 30 days. The pH dropped below 7.6 on June 17. It is now holding around 7.73.

screen shot 2017-06-30 at 10 54 31 am

Figure 6. AVTECH temperature readings for the past 10 days. The temperature held around 18ºC until June 25. Temperature reached a peak of 20ºC on June 26, dropping after an increase in flow. Temperature is now at 15.43ºC.

Now that I’ve stabilized pH, my next step is to revise my modified conditioning plan.

Written on June 30, 2017